# SECURA: Sigmoid-Enhanced CUR Decomposition with Uninterrupted Retention and Low-Rank Adaptation in Large Language Models

### v3

@@ 4.2 Main Result
As shown in table, most of PEFT variants almost lose all of its former knowledge, even the variants which aimed to relive catastrophic forgetting such as CUR-LoRA and I-LoRA are still have heavily forgetting and only can keep average score under 42.59%, while SECURA shows strong capabilities of limit catastrophic forgetting even in such tense environment and gains average score of 72.21% under the test of F-task. Demonstrating its robustness in mitigating catastrophic forgetting.
@@ 4.3 Ablation Study
Table 3: Ablation study comparing performance of 1-step and 200-step merges. The 1-step merge achieves better performance with minimal forgetting, while the 200-step merge leads to a cycle with reduced performance
@@ 5.2.2 No Need for Former Datasets and Calculation
5.2.3 Easy to Establish
Both EWC and Experience Replay require additional preparation steps before training begin, which adds complexity. In contrast, SECURA allows fine-tuning the model without any additional operations—just like fine-tuning a LoRA model. This makes SECURA much easier to implement and more accessible for practical applications.
@@ 6 Conclusion
We propose SECURA, a novel PEFT fine-tune method that integrates SigNorm normalization and CABR decomposition to address catastrophic forgetting in LLM fine-tuning. SECURA dynamically balances parameter updates using a Sigmoid-based pruning mechanism, preserving critical knowledge while adapting to new tasks. Experiments on 18 datasets and 5 LLMs show SECURA outperforms standard LoRA, with a 3.63% improvement on MCQ tasks and 2.56% on QA tasks, while retaining over 70% of base knowledge in continual learning scenarios. Paving the way for sustainable and ethical deployment of large AI models.
@@ 7 Limitations
  • Computational Overhead: SigNorm introduces additional matrix operations during training, increasing per-step time by approximately 1.18% compared to LoRA. Future work could focus on optimizing the normalization layer for real-time applications.
@@ Appendix A Broader Impacts
  • Generalization Beyond NLP: The SigNorm algorithm, though initially designed for PEFT, could be extended to other continual learning scenarios (e.g., robotics, healthcare), potentially improving adaptive systems in dynamic environments.
Input: Base weights $\mathcal{W}_{\text{former}}$, hyperparameters $r,m,\epsilon,\text{Scale}$
@@
1. Initialize CABR matrices:  $\mathcal{W}_{\text{frozen}}=USV^{T}$
@@
$\mathcal{W}_{A}\leftarrow U[:r,:]\cdot S\cdot V[m:,:]^{T}$
@@
Step 2: SigNorm Normalization
@@ 2. Merge weights:
$\mathcal{W}_{\text{SECURAMerged}}\leftarrow\mathcal{W}_{\text{CABR}}+\mathcal{W}_{\text{former}}$
@@ 3. Compute magnitude ratio:
$\mathcal{M}_{\text{Mag}}\leftarrow\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{former}}+\epsilon}\right|$
@@ 6. Adjust updates:
        $\mathcal{W}_{\text{Newformer}}\leftarrow\mathcal{W}_{\text{CABR}}+\mathcal{W}_{\text{former}}$
@@
        $\mathcal{W}_{\text{former}}\leftarrow CA_{\text{former}}B_{\text{former}}R+\mathcal{W}_{\text{basic}}$
        $\mathcal{W}_{\text{Newformer}}\leftarrow\mathcal{W}_{\text{former}}+\mathcal{W}_{\text{CABR}}$
@@ Appendix D Experimental hyper parameters on three main experiments
Table  6 and  7 show all hyper-parameters among different LoRA baselines and SECURA using in three main experiments.
@@
Table 6: hyper-parameters of all experiments
@@
CABR/CUR r Rank - - - - 150 -
CABR m Rank - - - - 200 -
Fusion Step 2000 2000 2000 2000 1 1
@@
Table 7: Hyper-parameters of all experiments
@@ Appendix E Experimental Results on All Datasets
Table 8910,  12 and  11 show all results among different LoRA baselines and SECURA using Gemma2 2b, Qwen2 1.5B, Qwen2 7B, LLaMA3 7B and LLaMA3.1 8B backbones.

### v4

@@ 4.2 Main Result
As shown in table, most of PEFT variants almost lose all of its former knowledge, even the variants which aimed to alleviate catastrophic forgetting such as CUR-LoRA and I-LoRA are still have heavily forgetting and only can keep average score under 42.59%, while SECURA shows strong capabilities of limit catastrophic forgetting even in such tense environment and gains average score of 72.21% under the test of F-task. Demonstrating its robustness in mitigating catastrophic forgetting.
@@ 4.3 Ablation Study
We have already conducted the ablation study comparing between Gradient variations and SVD norm in figure  2 and  3. Here, we provide a more detailed ablation study focusing on the issues arising from different merge methods.
@@
Table 3: Ablation study comparing knowledge retention performance of 1-step and 200-step merges under EXP2. The 1-step merge achieves better performance with minimal forgetting, while the 200-step merge leads to a cycle with reduced performance
@@ 6 Conclusion
We propose SECURA, a novel PEFT fine-tune method that integrates our newly proposed S-MagNorm normalization and CABR decomposition to address catastrophic forgetting in LLM fine-tuning. SECURA dynamically balances parameter updates using a Sigmoid-based pruning mechanism, preserving critical knowledge while adapting to new tasks. Experiments on 18 datasets and 5 LLMs show SECURA outperforms standard LoRA, with a 3.63% improvement on MCQ tasks and 2.56% on QA tasks, while retaining over 70% of base knowledge in continual learning scenarios. Paving the way for sustainable and ethical deployment of large AI models.
@@ 7 Limitations
  • Computational Overhead: S-MagNorm introduces additional matrix operations during training, increasing per-step time by approximately 1.18% compared to LoRA. Future work could focus on optimizing the normalization layer for real-time applications.
@@ Appendix A Broader Impacts
  • Generalization Beyond NLP: The S-MagNorm algorithm, though initially designed for PEFT, could be extended to other continual learning scenarios (e.g., robotics, healthcare), potentially improving adaptive systems in dynamic environments.
Input: Base weights $\mathcal{W}_{\text{base}}$, hyperparameters $r,m,\epsilon,\text{Scale}$
@@
1. Initialize CABR matrices:  $\mathcal{W}_{\text{frozen}}=\mathcal{U}_{\text{SVD}}\mathcal{S}_{\text{SVD}}\mathcal{V}_{\text{SVD}}^{\top}.$
@@
$\mathcal{W}_{A}\leftarrow\mathcal{U}_{\text{SVD}}^{(r,:)}\mathcal{S}_{\text{SVD}}\mathcal{V}_{\text{SVD}}^{(:,m)\top}$
@@
Step 2: S-MagNorm Normalization
@@ 2. Merge weights:
$\mathcal{W}_{\text{SECURAMerged}}\leftarrow\mathcal{W}_{\text{CABR}}+\mathcal{W}_{\text{base}}$
@@ 3. Compute magnitude ratio:
$\mathcal{M}_{\text{Mag}}\leftarrow\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{base}}+\epsilon}\right|$
@@ 6. Adjust updates:
        $\mathcal{W}_{\text{updated}}\leftarrow CA_{\text{former}}B_{\text{train}}R+\mathcal{W}_{\text{base}}$
@@
        $\mathcal{W}_{\text{mergedBase}}\leftarrow CA_{\text{frozen}}B_{\text{accum}}R+\mathcal{W}_{\text{base}}$
        $\mathcal{W}_{\text{merged}}\leftarrow\mathcal{W}_{\text{mergedBase}}+CA_{\text{train}}B_{\text{reset}}R$
@@ Appendix D Experimental hyper parameters on three main experiments
All of our main experiments conducted on two NVIDIA 2080 Ti 11GB GPUs, table  6,  7,  8 and  9 show all hyper-parameters among different LoRA baselines and SECURA using in experiments.
@@
Table 6: hyper-parameters of experiment 1&2
@@
CABR/CUR r Rank - - - - 150 150
CABR m Rank - - - - 200 200
Fusion Step 2000 2000 2000 1 2000 1
@@
Table 7: Hyper-parameters of experiment 3
Gradient Exp LoRA CABR-LoRA SECURA
Rank 8 - -
Learning Rates 1e-3 1e-3 1e-3
CABR/CUR r Rank - 150 150
CABR m Rank - 200 200
Fusion Step - - 1
Trainable Parameters (each layer) 2.62e+5 1.20e+5 1.00e+5
Table 8: Hyper-parameters of Gradient comparison experiments
SVD Norm Exp LoRA DoRA CABR-LoRA SECURA
Rank 8 8 - -
Learning Rates 1e-3 1e-3 1e-3 1e-3
CABR/CUR r Rank - - 150 150
CABR m Rank - - 200 200
Fusion Step - - - 1
Trainable Parameters (each layer) 2.62e+5 1.48e+5 1.20e+5 1.00e+5
Table 9: Hyper-parameters of SVD Norm experiments
@@ Appendix E Experimental Results on All Datasets
Table 101112,  14 and  13 show all results among different LoRA baselines and SECURA using Gemma2 2b, Qwen2 1.5B, Qwen2 7B, LLaMA3 7B and LLaMA3.1 8B backbones.